OpenStudy (anonymous):

Find an equation of the curve that passes through the point (0, 1) and whose slope at (x, y) is 5xy.

OpenStudy (dumbcow):

slope = dy/dx \[\frac{dy}{dx} = 5xy\] \[\frac{dy}{y}= 5x dx\] integrate both sides \[\ln y = (5/2) x^{2} +C\] \[y = k e^{2.5x^{2}}\] (k = e^C) plug in given point to find "k"